# Statistical Inference for Explainable Boosting Machines

### v1

@@ 5 Inference and Computation
  • $c_{E}^{-1}\widehat{\mathbf{f}}^{(k)}_{E}(\mathbf{x})\ \pm\ z_{1-\alpha/2}\,c_{E}^{-1}\hat{\sigma}\,\lVert\mathbf{r}^{(k)}_{E}(\mathbf{x})\rVert$ yields a $(1-\alpha)$-confidence interval for the feature effect. By similar arguments in (Fang et al., 2025), the coverage rate will converge to $(1-\alpha)$ almost surely. This is validated by an example in Figure 3.
  • For the intercept term, by Lemma E.15, $\widehat{\boldsymbol{\beta}}$ is asymptotically normal with variance $\sigma^{2}/n$, yielding the usual interval $\widehat{\boldsymbol{\beta}}\pm z_{1-\alpha/2}\,\hat{\sigma}/\sqrt{n}$
@@
(Fang et al., 2025) examines two further variants that are immediately implementable:
@@
Binning
The current implementation of EBMs (Nori et al., 2019) utilizes histogram trees. As we worked off this implementation, it was natural to construct intervals with the binned data, instead of working with the original data.
As such, we compress all $n$ samples along feature $k$ into $m_{k}\!\ll\!n$ histogram bins and work entirely in bin space. In Appendix C, we construct a decomposition of the structure matrix that allows us to show that computing the norm of $\mathbf{r}^{(k)}(\mathbf{x})$ reduces to solving a $m_{k}\times m_{k}$ system of linear equations, and computing two $m_{k}$-length dot products.
After this compression, the dependence on $n$ disappears. Denote $m$ as the maximum number of bins among all $k$ binnings. Caching $\mathbf{M}$ takes $O(pm^{3})$, while each per-point inference query takes only $O(m^{2})$ time. As bin numbers are often capped to 255 or 511 (independent of $n$), we can compute our intervals in time independent of the number of samples. This outperforms (Fang et al., 2025)’s runtime of $O(nd_{\text{eff}}^{2})$.
@@ 6 Numerical Experiments
Qualitatively, our intervals are better-behaved and much faster to compute than those of (Fang et al., 2025). This is no surprise, as we avoid the curse of dimensionality and our intervals have runtime independent of the number of samples.
Figure 5: RMSE for Algorithm 1 and benchmarks with 2 s.e. bands over 50 trials, on $f(x)=\sin(4\pi x)$ and the diabetes dataset from (Efron et al., 2004). Both EBM and Algorithm 1 are highly resistant to overfitting.
@@ 7 Conclusion
This work opens several extensions. We can readily extend the function class to more complex additive structure; including isolated interaction terms is straightforward, although additive components with shared features require new identifiability conditions (e.g. Lengerich et al., 2020). More generally, these methods are extensible to mixtures of learners, such as the class of varying coefficient models discussed in Zhou and Hooker (2022). We can an also broaden the class of models to discrete outcomes and models with non-constant variance. Finally, the kernel form of the limit also admits combinations and comparisons across models as described in (Ghosal et al., 2022).
@@
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes; See Sections 2 and 3, and Appendices C and B ]
@@
      Complete proofs of all theoretical results. [Yes; see Appendices D, E, and F]
@@
Appendix A Additional Literature

### v2

@@ 5 Inference and Computation
  • $c_{E}^{-1}\widehat{\mathbf{f}}^{(k)}_{E}(\mathbf{x})\ \pm\ z_{1-\alpha/2}\,c_{E}^{-1}\hat{\sigma}\,\lVert\mathbf{r}^{(k)}_{E}(\mathbf{x})\rVert$ yields a $(1-\alpha)$-confidence interval for the feature effect. By similar arguments in Fang et al. (2025), the coverage rate will converge to $(1-\alpha)$ almost surely. This is validated by an example in Figure 3.
  • For the intercept term, by Lemma F.15, $\widehat{\boldsymbol{\beta}}$ is asymptotically normal with variance $\sigma^{2}/n$, yielding the usual interval $\widehat{\boldsymbol{\beta}}\pm z_{1-\alpha/2}\,\hat{\sigma}/\sqrt{n}$
@@
Fang et al. (2025) examines two further variants that are immediately implementable:
@@
Binning.
The current implementation of EBMs Nori et al. (2019) utilizes histogram trees. As we worked off this implementation, it was natural to construct intervals with the binned data, instead of working with the original data. A more detailed discussion of the discretization error induced by binning is deferred to Appendix D.3.
As such, we compress all $n$ samples along feature $k$ into $m_{k}\!\ll\!n$ histogram bins and work entirely in bin space. In Appendix D, we construct a decomposition of the structure matrix that allows us to show that computing the norm of $\mathbf{r}^{(k)}(\mathbf{x})$ reduces to solving a $m_{k}\times m_{k}$ system of linear equations, and computing two $m_{k}$-length dot products.
After this compression, the dependence on $n$ disappears. Denote $m$ as the maximum number of bins among all $k$ binnings. Caching $\mathbf{M}$ takes $O(pm^{3})$, while each per-point inference query takes only $O(m^{2})$ time. As bin numbers are often capped to 255 or 511 (independent of $n$), we can compute our intervals in time independent of the number of samples. This outperforms Fang et al. (2025)’s runtime of $O(nd_{\text{eff}}^{2})$.
Practical Guidance.
In practice, Algorithm 1 relies on conditional orthogonality; we recommend monitoring simple convergence diagnostics, and if convergence appears unstable, falling back to Algorithm 2. In our experiments, the number of boosting rounds $B$ in the range 100–300 worked well. A smaller learning rate $\lambda$ typically improves finite-sample stability but can yield slightly wider intervals. Subsample rates in $[0.5,0.8]$ were robust. For binning, we found that using either the usual EBM defaults (e.g., 64 or 255 bins) or scaling the bin count as $m\asymp n^{1/3}$ works well; the latter matches our assumptions.
@@ 6 Numerical Experiments
We further evaluate scalability in higher-dimensional, correlated synthetic additive models for Algorithm 2 under the realistic Assumption 4.9. In Figure 1, bootstrap EBMs achieve low RMSE but severely under-cover and are slow due to repeated refits, while Algorithm 2 (and BRATD) attain near-nominal coverage; moreover, Algorithm 2 yields better RMSE and tighter intervals than BRATD at comparable or lower runtime. Compared with Fang et al. (2025), our intervals are also faster and better-behaved, since we avoid the curse of dimensionality and our runtime is independent of $n$.
Figure 2 shows that on real datasets with $p\approx 110$$130$, Algorithm 2 maintains near-nominal $95\%$ coverage (0.9480.9480.9560.956), achieves RMSE comparable to the baselines, and is substantially faster than vanilla EBM. Relative to BRATD, it attains similar RMSE with lower training time on three of the four datasets, especially on the largest example.
Figure 4: RMSE for Algorithm 1 and benchmarks with 2 s.e. bands over 50 trials, on $f(x)=\sin(4\pi x)$ and the diabetes dataset from Efron et al. (2004). Both EBM and Algorithm 1 are highly resistant to overfitting.
@@ 7 Conclusion
This work opens several extensions. We can readily extend the function class to more complex additive structure; including isolated interaction terms is straightforward, although additive components with shared features require new identifiability conditions (e.g. Lengerich et al., 2020). More generally, these methods are extensible to mixtures of learners, such as the class of varying coefficient models discussed in Zhou and Hooker (2022). We can an also broaden the class of models to discrete outcomes and models with non-constant variance. Finally, the kernel form of the limit also admits combinations and comparisons across models as described in Ghosal et al. (2022).
$p$ Method RMSE Coverage Width Time (s)
5 Bootstrap EBM 0.514 0.290 0.386 26.42
Algorithm 2 0.661 0.951 2.609 3.00
BRATD 0.611 0.956 2.899 10.41
10 Bootstrap EBM 0.537 0.373 0.545 21.24
Algorithm 2 0.788 0.951 3.069 5.90
BRATD 0.866 0.935 3.531 13.58
15 Bootstrap EBM 0.563 0.453 0.675 31.51
Algorithm 2 0.900 0.945 3.520 8.77
BRATD 1.112 0.946 4.398 10.29
20 Bootstrap EBM 0.583 0.511 0.801 43.14
Algorithm 2 1.004 0.948 3.901 11.56
BRATD 1.332 0.948 5.121 10.70
Table 1: Synthetic Lipschitz GAMs (bootstrap EBM vs. Algorithm 2 vs. BRATD). Synthetic additive model with $n=1000$ and $p\in\{5,10,15,20\}$ (all $p$ features are structural; five base univariate functions are cycled across coordinates): $f_{1}(x)=2\sin(\pi x)$, $f_{2}(x)=(x-0.5)^{2}$, $f_{3}(x)=0.5\times\mathbf{1}\{x>0.3\}$, $f_{4}(x)=\sqrt{x}$, $f_{5}(x)=-1.5x$. We report mean RMSE, model-level empirical coverage at nominal $95\%$, average interval width, and runtime (seconds) over 10 repetitions ($\alpha=0.05$; 250 bootstrap resamples for EBM).
Dataset $n$ $p$
RMSE
(Alg. 2)
Cov.
(Alg. 2)
Time (s)
(Alg. 2)
taiwanese_bankruptcy 6819 95 0.151 0.948 18.01
myocardial_infarction 1700 111 0.309 0.943 3.53
communities_and_crime 1994 127 0.140 0.953 5.68
dota2_results 50000 116 0.142 0.956 28.78
Dataset
RMSE
Time (s)
RMSE
Time (s)
taiwanese_bankruptcy 0.153 197.31 0.154 24.58
myocardial_infarction 0.309 52.62 0.311 3.78
communities_and_crime 0.135 178.41 0.144 7.80
dota2_results 0.013 3736.10 0.147 82.62
Table 2: We report RMSE, $95\%$ empirical coverage, and training time for Algorithm 2 (top), and baseline RMSE and training time for vanilla EBM and BRATD (bottom). Baseline coverage is omitted because large-scale uncertainty estimation is not computationally scalable.
@@
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes; See Sections 2 and 3, and Appendices D and C ]
@@
      Complete proofs of all theoretical results. [Yes; see Appendices E, F, and G]
@@
Appendix A Notations
Symbol Meaning / definition (with dimensions when helpful)
$n$ Sample size; data $(\mathbf{X}_{n},\mathbf{y}_{n})=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$ are i.i.d.
$p$ Number of features (coordinates) in $\mathbf{x}\in\mathbb{R}^{p}$.
$i,j,\ell$ Generic indices: $i,j\in\{1,\dots,n\}$ for samples; $\ell\in\{1,\dots,p\}$ for features.
$\mathbf{x}_{i}$ $i$-th covariate vector: $\mathbf{x}_{i}=(\mathbf{x}_{i}^{(1)},\dots,\mathbf{x}_{i}^{(p)})^{\top}\in\mathbb{R}^{p}$.
$\mathbf{x}_{i}^{(k)}$ Feature-$k$ coordinate of $\mathbf{x}_{i}$ (scalar in the univariate EBM setting).
$\mathbf{X}\in\mathbb{R}^{n\times p}$ Design matrix with rows $\mathbf{x}_{i}^{\top}$.
$y_{i}$ Response for sample $i$ (scalar).
$\mathbf{y}\in\mathbb{R}^{n}$ Response vector $(y_{1},\dots,y_{n})^{\top}$.
$\epsilon$ / $\varepsilon$ Noise; i.i.d. sub-Gaussian errors with $\mathbb{E}[\varepsilon]=0$, $\text{Var}(\varepsilon)=\sigma^{2}<\infty$.
$\sigma^{2}$ Noise variance in Assumption 1.
$\mu(\mathbf{x})$ Density of covariates on their support; bounded as $0<c_{1}\leq\mu(\mathbf{x})\leq c_{2}<\infty$.
$\boldsymbol{\beta}$ True intercept: $\boldsymbol{\beta}:=\mathbb{E}_{\mathbf{x},\epsilon}[y]$ (unconditional mean of $y$).
$\widehat{\boldsymbol{\beta}}_{b}$ Estimated intercept at boosting round $b$ (Algorithm 1 / Alg. in appendix). Initialized as $\widehat{\boldsymbol{\beta}}_{0}=\frac{1}{n}\mathbf{1}^{\top}\mathbf{y}$.
$\bar{\mathbf{y}}$ Sample mean of responses: $\bar{\mathbf{y}}:=\frac{1}{n}\mathbf{1}^{\top}\mathbf{y}$.
$\mathbf{f}(\mathbf{x})$ True regression function in GAM form: $\mathbf{f}(\mathbf{x})=\boldsymbol{\beta}+\sum_{k=1}^{p}\mathbf{f}^{(k)}(\mathbf{x}^{(k)})$.
$\mathbf{f}^{(k)}$ True univariate component function for feature $k$; $L$-Lipschitz and mean-zero w.r.t. marginal of $\mathbf{x}^{(k)}$.
$\widehat{\mathbf{f}}$ Generic fitted predictor; in Algorithm 1 output is $\widehat{\mathbf{f}}(\mathbf{x})=\widehat{\boldsymbol{\beta}}_{B}+\frac{1+\lambda}{\lambda}\sum_{k=1}^{p}\mathbf{f}_{B}^{(k)}(\mathbf{x}^{(k)})$.
$\mathbf{f}_{b}^{(k)}$ Algorithm’s stored additive component for feature $k$ after round $b$ (a function of $\mathbf{x}^{(k)}$). Initialized $\mathbf{f}_{0}^{(k)}\equiv 0$.
$\widehat{\mathbf{y}}_{b}^{(k)}$ Feature-$k$ vector prediction on training points at round $b$ (used in Theorems): $\widehat{\mathbf{y}}_{b}^{(k)}\in\mathbb{R}^{n}$.
$\widehat{\mathbf{y}}_{b}$ Total fitted values on training set at round $b$: $\widehat{\mathbf{y}}_{b}:=\widehat{\boldsymbol{\beta}}_{b}+\sum_{a=1}^{p}\widehat{\mathbf{y}}_{b}^{(a)}$ (Theorem 1).
$\mathbf{1}$ All-ones vector in $\mathbb{R}^{n}$.
$\mathbf{I}_{n}$ / $\mathbf{I}$ Identity matrix in $\mathbb{R}^{n\times n}$ (often written $\mathbf{I}$ when $n$ is clear).
$\mathbf{J}_{n}$ / $\mathbf{J}_{n}$ Centering projector: $\mathbf{J}_{n}=\mathbf{I}_{n}-\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}$.
Table 3: Basic data & model notation.
Symbol Meaning / definition (with dimensions when helpful)
$b$ Boosting round index ($b=1,2,\dots$).
$B$ Number of boosting rounds (loop bound in pseudocode).
$\lambda$ Learning rate in Boulevard-style moving average update.
$\xi$ Subsample rate: each $G_{b,k}\subset\{1,\dots,n\}$ includes each index i.i.d. with prob. $\xi$.
$M$ Truncation / clipping level: $\widetilde{t}\leftarrow\max\{-M,\min\{\widetilde{t},M\}\}$.
$G_{b,k}$ Subsample index set used to fit tree for feature $k$ at round $b$.
$r_{i,k}$ Residual used to fit feature-$k$ tree at round $b$. Alg 1: $r_{i,k}=y_{i}-\big(\widehat{\boldsymbol{\beta}}_{b-1}+\sum_{\ell}\mathbf{f}_{b-1}^{(\ell)}(\mathbf{x}_{i}^{(\ell)})\big)$. Alg 2 (leave-one-out): sum excludes $\ell=k$.
$\mathbf{t}^{(b,k)}$ Regression tree trained at round $b$ on feature $k$ using pairs $\{(\mathbf{x}_{i}^{(k)},r_{i,k})\}_{i\in G_{b,k}}$.
$\mu_{b,k}$ Mean (centering constant) of tree predictions on training set: $\mu_{b,k}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{t}^{(b,k)}(\mathbf{x}_{i}^{(k)})$.
$\widetilde{\mathbf{t}}_{b}^{(k)}$ Centered (and then clipped) tree: $\widetilde{\mathbf{t}}_{b}^{(k)}(\mathbf{x})=\mathbf{t}^{(b,k)}(\mathbf{x})-\mu_{b,k}$, then clipped to [M,M][-M,M].
Table 4: Algorithmic hyperparameters & indices.
Symbol Meaning / definition (with dimensions when helpful)
$\mathbf{t}_{n}$ A regression tree trained on $n$ datapoints; induces a partition $\{A_{i}\}_{i=1}^{m}$ of $0,1]^{d}$ (Def. 1).
$A(\mathbf{x})$ The leaf (rectangle) containing $\mathbf{x}$ under the tree partition.
$s_{n,j}(\mathbf{x})$ Structure weight of training point $j$ for query $\mathbf{x}$: $s_{n,j}(\mathbf{x})=\frac{\mathbbm{1}(\mathbf{x}_{j}\in A(\mathbf{x}))}{\sum_{k=1}^{n}\mathbbm{1}(\mathbf{x}_{k}\in A(\mathbf{x}))}$.
$\mathbf{s}_{n}(\mathbf{x})$ Structure vector: $\mathbf{s}_{n}(\mathbf{x})=(s_{n,1}(\mathbf{x}),\dots,s_{n,n}(\mathbf{x}))^{\top}\in\mathbb{R}^{n}$.
$\mathbf{S}_{n}$ Structure matrix: $\mathbf{S}_{n}=(s_{n,j}(\mathbf{x}_{i}))_{i,j=1}^{n}\in\mathbb{R}^{n\times n}$; row $i$ is $\mathbf{s}_{n}(\mathbf{x}_{i})^{\top}$.
$\mathbf{S}^{(k)}_{b}$ Feature-$k$ structure matrix at boosting round $b$ (Assumption 2). [I interpret $\mathbf{S}^{(k)}_{b}$ as $\mathbf{S}_{n}$ for the univariate tree fit on feature $k$ at round $b$.]
$\mathbf{K}^{(k)}$ Expected feature-$k$ kernel: $\mathbf{K}^{(k)}:=\mathbb{E}[\mathbf{S}^{(k)}]$ (also written $\mathbb{E}[\mathbf{S}^{(k)}]$).
$\mathbf{K}$ Aggregated kernel (main text / Thm 1): $\mathbf{K}:=\sum_{k=1}^{p}\mathbf{K}^{(k)}$.
$\mathcal{Q}_{n}^{(k)}$ Distribution over feature-$k$ tree structures after burn-in (Assumption 2).
$Q_{n}^{(k)}$ Tree-structure support / “tree space” for feature $k$ (Assumptions 3–5).
$b^{\prime}$ Burn-in iteration after which structures are i.i.d. from $\mathcal{Q}_{n}^{(k)}$ (Assumption 2).
Table 5: Structure vectors & matrices.
Symbol Meaning / definition (with dimensions when helpful)
$\Pi$ A tree partition (set of leaves) drawn from $Q_{n}^{(k)}$.
$\mathsf{diam}(A)$ Leaf diameter: $\sup_{\mathbf{x}_{1},\mathbf{x}_{2}\in A}\|\mathbf{x}_{1}-\mathbf{x}_{2}\|$.
$d_{n}$ Target diameter scale: $d_{n}=O(n^{-1/3})=o(1/\log n)$ (Assumption 3).
$v_{n}$ Leaf geometric volume bound, with $v_{n}=n^{-2/3+\nu}$ for any $\nu>0$ (Assumption 4).
$\epsilon_{n}$ Small sequence $\epsilon_{n}\to 0$ in restricted tree support bound (Assumption 5).
Table 6: Geometric & complexity parameters for trees.
Symbol Meaning / definition (with dimensions when helpful)
$\widetilde{\mathbf{y}}_{k}^{*}$ Feature-$k$ fixed point vector (Thm 1 / Thm 2). In Thm 1: $\widetilde{\mathbf{y}}_{k}^{*}=\mathbf{J}_{n}\,\mathbf{K}^{(k)}[\lambda^{-1}\mathbf{I}+\mathbf{J}_{n}\mathbf{K}]^{-1}\mathbf{y}$.
$\widehat{\boldsymbol{\beta}}^{*}$ Intercept fixed point: $\widehat{\boldsymbol{\beta}}^{*}=\bar{\mathbf{y}}$.
$\widehat{\mathbf{y}}^{*}$ Total fixed point fit on training set: $\widehat{\mathbf{y}}^{*}=\widehat{\boldsymbol{\beta}}^{*}\mathbf{1}+\sum_{k=1}^{p}\widetilde{\mathbf{y}}_{k}^{*}$.
$\mathcal{K}$ “combined kernel” for Algorithm 2: $\mathcal{K}:=\sum_{k=1}^{p}(\mathbf{I}-\mathbb{E}\mathbf{S}^{(k)})^{-1}\mathbf{J}_{n}\mathbb{E}\mathbf{S}^{(k)}$.
$\mathcal{K}_{\text{bin}}$ $\mathcal{K}_{\text{bin}}=\sum_{k=1}^{p}\mathbb{E}[\mathbf{b}_{n}^{(k)}\mathbf{B}^{(k)}\mathbf{b}_{n}^{(k)\top}]$.
$\mathbf{r}^{(k)}$ / $\mathbf{r}^{(k)}_{E}(\mathbf{x})$ Kernel-ridge “influence / weight” vector used in CLTs and intervals; satisfies (example for Alg 1): ${\mathbf{r}^{(k)}}^{\top}\mathbf{y}=\mathbb{E}[\mathbf{s}^{(k)}(\mathbf{x})]^{\top}\mathbf{J}_{n}[\mathbf{I}+\mathbf{J}_{n}\mathbf{K}]^{-1}\mathbf{y}$. Subscript $E\in\{A,B\}$ labels algorithm family (Alg 1 vs Alg 2) in Theorem 3.
$\mathbf{k}^{(k)}$ / $\mathbf{k}^{(k)}_{n}(\mathbf{x})$ Kernel vector (expected structure vector / feature-$k$ similarity to training points). In binning appendix: $\mathbf{k}^{(k)}_{n}(\mathbf{x})=\mathbb{E}[\mathbf{b}^{(k)}(\mathbf{x})^{\top}\mathbf{D}^{(k)}\mathbf{b}_{n}^{(k)\top}]\in\mathbb{R}^{1\times n}$.
$c_{E}$ Algorithm-dependent scaling constant in CLT: $c_{A}=\lambda/(1+\lambda)$ for Algorithms 1, 3; $c_{B}=1$ for Algorithm 2 (Theorem 3).
$z_{1-\alpha/2}$ Standard normal quantile used for $(1-\alpha)$ intervals.
Table 7: Fixed points, CLT scaling, weight vectors.
Symbol Meaning / definition (with dimensions when helpful)
$\mathcal{H}_{k}$ Centered feature-$k$ additive subspace: $\mathcal{H}_{k}=\{\mathbf{v}\in\mathbb{R}^{n}:\exists g:\mathbb{R}\to\mathbb{R},\ \mathbf{v}_{i}=g(\mathbf{x}_{i}^{(k)}),\ \mathbf{1}^{\top}\mathbf{v}=0\}$.
$P_{k}$ Orthogonal projector onto $\mathcal{H}_{k}$.
$\varepsilon_{n,k}$ Operator-norm approximation error s.t. $\|\mathbf{J}_{n}\mathbf{K}^{(k)}\mathbf{J}_{n}-P_{k}\|_{\mathrm{op}}\leq\varepsilon_{n,k}\to 0$.
Table 8: Additive subspaces and projectors; Assumption 4.7).
Symbol Meaning / definition (with dimensions when helpful)
$m_{k}$ Number of histogram bins for feature $k$; $m:=\max_{k}m_{k}$.
$B_{r}^{(k)}$ The $r$-th bin for feature $k$ (a subset/interval of $\mathbb{R}$).
$r_{i}^{(k)}$ Bin index of sample $i$ on feature $k$: unique $r$ with $\mathbf{x}_{i}^{(k)}\in B_{r}^{(k)}$.
$\mathcal{L}^{(k)}(r)$ Set of bin indices in the same tree leaf as bin $r$ (feature $k$): $\{s\in\{1,\dots,m_{k}\}:B_{s}^{(k)}\in A(B_{r}^{(k)})\}$.
$n_{s}^{(k)}$ Bin count: number of samples in bin $s$ for feature $k$, $n_{s}^{(k)}=\sum_{j=1}^{n}\mathbf{1}(\mathbf{x}_{j}^{(k)}\in B_{s}^{(k)})$.
$N_{\mathcal{L}(r)}^{(k)}$ Leaf sample count for the leaf containing bin $r$: $N_{\mathcal{L}(r)}^{(k)}=\sum_{q\in\mathcal{L}^{(k)}(r)}n_{q}^{(k)}$.
$|\mathcal{L}^{(k)}(r)|$ Leaf bin count (number of bins in that leaf).
$\mathbf{Z}_{n}^{(k)}\in\{0,1\}^{n\times m_{k}}$ Sample-to-bin assignment matrix: $\mathbf{Z}_{i,j}^{(k)}=\mathbf{1}(r_{i}^{(k)}=j)=\mathbf{1}(\mathbf{x}_{i}^{(k)}\in B_{j}^{(k)})$.
$\mathbf{D}^{(k)}=\mathrm{diag}(\xi_{1}^{(k)},\dots,\xi_{m_{k}}^{(k)})$ Diagonal rescaling to convert bin-normalization to sample-normalization, with $\xi_{r}^{(k)}=\sqrt{\frac{|\mathcal{L}^{(k)}(r)|}{N_{\mathcal{L}(r)}^{(k)}}}$.
$\mathbf{B}^{(k)}\in\mathbb{R}^{m_{k}\times m_{k}}$ Bin-structure matrix: $\mathbf{B}^{(k)}_{i,j}=\frac{\mathbf{1}(B_{j}^{(k)}\in A(B_{i}^{(k)}))}{\sum_{l=1}^{m_{k}}\mathbf{1}(B_{l}^{(k)}\in A(B_{i}^{(k)}))}$.
$\mathbf{b}_{n}^{(k)}$ Matrix mapping bin-space to sample-space. Initially given as $\mathbf{b}_{n}^{(k)}\in\mathbb{R}^{n\times m_{k}}$; later simplified to $\mathbf{b}_{n}^{(k)}=\mathbf{Z}_{n}^{(k)}\mathbf{D}^{(k)}$.
$\mathbf{b}^{(k)}(\mathbf{x})\in\mathbb{R}^{m_{k}}$ Bin-space structure vector for query $\mathbf{x}$: $\mathbf{b}^{(k)}(\mathbf{x})^{\top}=\left(\frac{\mathbf{1}(B_{j}^{(k)}\in A(\mathbf{x}))}{\sum_{l=1}^{m_{k}}\mathbf{1}(B_{l}^{(k)}\in A(\mathbf{x}))}\right)_{j=1}^{m_{k}}$.
$\mathbf{h}^{(k)}(\mathbf{x})$ Defined in appendix lemma: $\mathbf{h}^{(k)}(\mathbf{x})=\mathbb{E}[(\mathbf{D}^{(k)})^{2}\mathbf{b}^{(k)}(\mathbf{x})]\in\mathbb{R}^{m_{k}}$.
$\mathbf{H}$ Sum of expected bin-kernels: $\mathbf{H}=\sum_{k=1}^{p}\mathbb{E}[\mathbf{D}^{(k)}\mathbf{B}^{(k)}\mathbf{D}^{(k)}]$.
$\mathbf{c}^{(k)}$ Bin counts vector: $\mathbf{c}^{(k)}:=\mathbf{Z}_{n}^{(k)\top}\mathbf{1}\in\mathbb{R}^{m_{k}}$.
Table 9: Binning & histogram-tree appendix.
Appendix B Additional Literature